What is the molarity of a solution if 200. mL of solution contains 5.00 g of Na2CO3?
1
Expert's answer
2012-05-09T10:58:07-0400
n=m/Mr
n=5.00/106=0.047 mol

Molarity is n/1L,so in 200 mL is 0.047 mol
and in 1000 mL (1 L)=0.188 mol
it is (0.188 mol/1000mL) 0.188 M